I'm in a pre-algebra class and we're working on solving linear equations with fractions and I'm not understanding it.
4/6 = x+8/15
how do I solve this?
Please help, thank you.

Answer by ReadingBoosters(3246)

Starting with
4/6 = x + 8/15
First isolate the x
x = 4/6 - 8/15
Then apply common denominators to your fractions
x = 4/6(5/5) - 8/15(2/2)
x = 20/30 - 16/30
Now you can subtract and simplify
x = 4/30 = 2/15
Post more if you have more questions and need further clarity.
.......................
Delighted to help.
-Reading Boosters
Wanting for others what we want for ourselves.